# A Covariant and Manifestly Projective Invariant Formulation of Thomas-Whitehead Gravity

### v1

@@ 3.3 Covariant and Manifestly Projective Invariant TW Field Equations
¯${\bar{\nabla}}_{a}A^{a}=$ $\partial_{a}A^{a}+{\bar{\Gamma}}^{a}{}_{ba}A^{a}$ (3.22)
$=$ $\partial_{a}A^{a}+\hat{\Gamma}^{a}{}_{ba}A^{a}+{\tilde{C}}^{a}{}_{ba}A^{a}$ (3.23)
$=$ $\partial_{a}A^{a}+\hat{\Gamma}^{a}{}_{ba}A^{a}$ (3.24)
$=$ ^$\hat{\nabla}_{a}A^{a}~{}~{}~{}.$ (3.25)
@@
$\int d^{\rm d}x\sqrt{g}\bar{\nabla}_{a}A^{a}=$ $\int d^{\rm d}x\sqrt{g}\hat{\nabla}_{a}A^{a}$ (3.26)
@@
$=$ $\int d^{\rm d}x\partial_{a}(\sqrt{g}A^{a})=\text{surface term}$ (3.29)
where we have used Eq. (2.30) in going from the second to third lines. This means we can use partial integration with ¯$\bar{\nabla}_{a}$ at the Lagrangian density level at no cost to the field equations. For instance given an arbitrary rank three tensor $T^{abc}$
$T^{abc}\bar{\nabla}_{a}\delta\bar{{\mathcal{P}}}_{bc}=$ (¯$-(\bar{\nabla}_{a}T^{abc})\delta\bar{{\mathcal{P}}}_{bc}$ (3.30)
up to a surface term. Thus, we can work entirely with the TW Lagrangian density
$\mathcal{L}=-\frac{1}{2\kappa_{0}}\mathcal{K}d^{\rm d}x+J_{0}c\lambda_{0}^{2}K_{bcd}K^{bcd}-J_{0}c\left(g_{am}\mathcal{K}^{a}_{\ bcd}\mathcal{K}^{mbcd}-4\mathcal{K}_{bd}\mathcal{K}^{bd}+\mathcal{K}^{2}\right)$ (3.31)
and need not write out the full action.
The connection ¯$\bar{\nabla}_{a}$ provides as much calculational functionality in TW gravity as the Levi-Civita connection ^$\hat{\nabla}_{a}$ does in conventional general relativity. This is a powerful new feature that the projectively invariant and tensorial ¯$\bar{\nabla}_{a}$ and ${\bar{{\mathcal{P}}}}_{ab}$ provide for TW gravity, moving the results of the previous work [20] closer to the language of more conventional language in general relativity. In the following, we will demonstrate the utility of this new calculation tool.
3.3.1 Field Equations for ${\bar{{\mathcal{P}}}}_{ab}$
@@
$\delta_{{\bar{{\mathcal{P}}}}}{\mathcal{K}}^{a}{}_{bcd}=$ $\delta^{a}{}_{[c}\delta{\bar{{\mathcal{P}}}}_{d]b}~{}~{}~{},~{}~{}~{}\delta_{{\bar{{\mathcal{P}}}}}K_{bcd}={\bar{\nabla}}_{[c}\delta{\bar{{\mathcal{P}}}}_{d]b}$ (3.32)
@@
$\delta_{{\bar{{\mathcal{P}}}}}\mathcal{L}=$ $-\tfrac{1}{2\kappa_{0}}g^{ab}\delta_{{\bar{{\mathcal{P}}}}}{\mathcal{K}}_{ab}+2J_{0}\lambda_{0}^{2}K^{bcd}\delta_{{\bar{{\mathcal{P}}}}}K_{bcd}$ (3.33)
$-2J_{0}\left(g_{am}{\mathcal{K}}^{mbcd}\delta_{{\bar{{\mathcal{P}}}}}{\mathcal{K}}^{a}{}_{bcd}+(-4{\mathcal{K}}^{ab}+{\mathcal{K}}g^{ab})\delta_{{\bar{{\mathcal{P}}}}}{\mathcal{K}}_{ab}\right)$ (3.34)
$=$ $-\tfrac{1}{2\kappa_{0}}g^{ab}({\rm d}-1)\delta{\bar{{\mathcal{P}}}}_{ab}+2J_{0}\lambda_{0}^{2}K^{bcd}{\bar{\nabla}}_{[c}\delta{\bar{{\mathcal{P}}}}_{d]b}$ (3.35)
$-2J_{0}(g_{am}{\mathcal{K}}^{mbcd}\delta^{a}{}_{[c}\delta{\bar{{\mathcal{P}}}}_{d]b}+(-4{\mathcal{K}}^{ab}+{\mathcal{K}}g^{ab})({\rm d}-1)\delta{\bar{{\mathcal{P}}}}_{ab})~{}~{}~{}.$ (3.36)
Integrating by parts in the second term, contracting indices in the third, and renaming and reordering indices yields
$\delta_{{\bar{{\mathcal{P}}}}}\mathcal{L}=$ $\big{(}-\tfrac{{\rm d}-1}{2\kappa_{0}}g^{ab}+2J_{0}\lambda_{0}^{2}{\bar{\nabla}}_{c}K^{(ab)c}$ (3.37)
$\hskip 20.0pt-2J_{0}({\rm d}-1){\mathcal{K}}g^{ab}+8J_{0}({\rm d}-1){\mathcal{K}}^{ab}-4J_{0}{\mathcal{K}}^{ab}\big{)}\delta{\bar{{\mathcal{P}}}}_{ab}$ (3.38)
where we have used the antisymmetry of the last two indices of $K^{abc}$ to reorder the indices. Simplifying we have
$\delta_{{\bar{{\mathcal{P}}}}}\mathcal{L}=$ $2J_{0}\mathbb{D}^{ab}\delta{\bar{{\mathcal{P}}}}_{ab}$ (3.39)
@@
$\mathbb{D}^{ab}=$ $-\frac{d-1}{4J_{0}c\kappa_{0}}g^{ab}-(d-1)g^{ab}\mathcal{K}+2(2d-3)\mathcal{K}^{ab}+\lambda_{0}^{2}\bar{\nabla}_{c}K^{(ab)c}=0~{}~{}~{}.$ (3.40)
Using Eq. (3.4) to expand the last term, it is straightforward to show that this is the exact same as Eq. (2.70) that was derived via variation with respect to ${\mathcal{D}}_{ab}$. In the form of Eq. (3.40), however, we see this equation of motion is manifestly covariant and projectively invariant as it is written entirely in terms of such objects.
3.3.2 Field Equations for ${\tilde{C}}^{c}{}_{ab}$
@@
$\delta_{{\bar{\Gamma}}}{\mathcal{K}}^{a}{}_{bcd}=$${\bar{\nabla}}_{[c}\delta{\bar{\Gamma}}^{a}{}_{d]b}~{}~{}~{},~{}~{}~{}\delta_{{\bar{\Gamma}}}K_{bcd}=-\delta{\bar{\Gamma}}^{m}{}_{b[c}{\bar{{\mathcal{P}}}}_{d]m}$ (3.41)
@@
$\delta_{{\bar{\Gamma}}}\mathcal{L}=$ $-\tfrac{1}{2\kappa_{0}}g^{ab}\delta_{{\bar{\Gamma}}}{\mathcal{K}}_{ab}+2J_{0}\lambda_{0}^{2}K^{bcd}\delta_{{\bar{\Gamma}}}K_{bcd}$ (3.42)
$-2J_{0}\left(g_{am}{\mathcal{K}}^{mbcd}\delta_{{\bar{\Gamma}}}{\mathcal{K}}^{a}{}_{bcd}+(-4{\mathcal{K}}^{ab}+{\mathcal{K}}g^{ab})\delta_{{\bar{\Gamma}}}{\mathcal{K}}_{ab}\right)$ (3.43)
$=$ $-\tfrac{1}{2\kappa_{0}}g^{ab}{\bar{\nabla}}_{[c}\delta{\bar{\Gamma}}^{c}{}_{b]a}+2J_{0}\lambda_{0}^{2}K^{bcd}(-\delta{\bar{\Gamma}}^{a}{}_{b[c}{\bar{{\mathcal{P}}}}_{d]m})$ (3.44)
$-2J_{0}(g_{am}{\mathcal{K}}^{mbcd}{\bar{\nabla}}_{[c}\delta{\bar{\Gamma}}^{a}{}_{d]b}+(-4{\mathcal{K}}^{ab}+{\mathcal{K}}g^{ab}){\bar{\nabla}}_{[c}\delta{\bar{\Gamma}}^{c}{}_{b]a})~{}~{}~{}.$ (3.45)
@@
Integrating by parts, relabeling some indices, and collecting like terms and simplifying results in
$\delta_{{\bar{\Gamma}}}\mathcal{L}=$ $\delta{\bar{\Gamma}}^{c}{}_{a[b}\left\{{\bar{\nabla}}_{c]}\left[\tfrac{1}{2\kappa_{0}}g^{ab}+2J_{0}\left(-4{\mathcal{K}}^{ab}+{\mathcal{K}}g^{ab}\right)\right]\right\}$ (3.46)
@@
$+\tfrac{1}{2}\left[-4J_{0}\lambda_{0}^{2}K^{(ab)d}{\bar{{\mathcal{P}}}}_{dc}+4J_{0}{\bar{\nabla}}_{d}\left(g_{cm}{\mathcal{K}}^{m(b|d|a)}\right)\right]\delta{\bar{\Gamma}}^{c}{}_{ab}~{}~{}~{}.$ (3.47)
@@
Introducing a Kronecker delta symbols in the first term to move the antisymmetrization of indices off of $\delta{\bar{\Gamma}}^{c}{}_{ab}$ and ¯${\bar{\nabla}}_{c}$, using the antisymmetry property ${\mathcal{K}}^{mbda}=-{\mathcal{K}}^{mbad}$ in the second term on the second line, and simplifying results in
$\delta_{{\bar{\Gamma}}}\mathcal{L}=$ $2J_{0}\bigg{[}{\bar{\nabla}}_{m}\left(\tfrac{1}{4J_{0}\kappa_{0}}g^{a[b}\delta_{c}{}^{m]}-4{\mathcal{K}}^{a[b}\delta_{c}{}^{m]}+{\mathcal{K}}g^{a[b}\delta_{c}{}^{m]}\right)$ (3.48)
$\hskip 70.0pt-\lambda_{0}^{2}{\bar{{\mathcal{P}}}}_{dc}K^{(ab)d}-{\bar{\nabla}}_{d}\left(g_{cm}{\mathcal{K}}^{m(ab)d}\right)\bigg{]}\delta{\bar{\Gamma}}^{c}{}_{ab}~{}~{}~{}.$ (3.49)
Defining
$\mathbb{E}_{c}{}^{ab}=$${\bar{\nabla}}_{c}\left[g^{ab}\left({\mathcal{K}}+\tfrac{1}{4J_{0}\kappa_{0}}\right)-4{\mathcal{K}}^{ab}\right]-{\bar{\nabla}}_{d}\left(g_{cm}{\mathcal{K}}^{m(ab)d}\right)-\lambda_{0}^{2}{\bar{{\mathcal{P}}}}_{dc}K^{(ab)d}$ (3.50)
the variation Eq. (3.48) can be cast in the following form
$\delta_{{\bar{\Gamma}}}\mathcal{L}=$ $2J_{0}\left[\mathbb{E}_{c}{}^{ab}-{\bar{\nabla}}_{m}\left(\tfrac{1}{4J_{0}\kappa_{0}}g^{am}-4{\mathcal{K}}^{am}+{\mathcal{K}}g^{am}\right)\delta_{c}{}^{b}\right]\delta{\bar{\Gamma}}^{c}{}_{ab}~{}~{}~{}.$ (3.51)
@@
To facilitate the goal of simultaneously isolating the equations of motion for ${\tilde{C}}^{c}{}_{ab}$ whilst keeping track of the contributions to the equations of motion for the metric $g^{mn}$, we define the traceless combination
$\tilde{\mathbb{E}}_{c}{}^{ab}\equiv$ $\mathbb{E}_{c}{}^{ab}-\tfrac{1}{{\rm d}+1}\delta_{c}{}^{(a}\mathbb{E}_{m}{}^{b)m}$ (3.52)
to express Eq. (3.51) as
@@
$\delta_{{\bar{\Gamma}}}\mathcal{L}=$ $2J_{0}\left\{\tilde{\mathbb{E}}_{c}{}^{ab}+\tfrac{1}{{\rm d}+1}\delta_{c}{}^{(a}\mathbb{E}_{m}{}^{b)m}k\right.$ (3.53)
$\qquad\left.-\tfrac{1}{2}{\bar{\nabla}}_{m}\left[\delta_{c}{}^{(a}g^{b)m}\left(\tfrac{1}{4J_{0}\kappa_{0}}+{\mathcal{K}}\right)-4\delta_{c}{}^{(a}{\mathcal{K}}^{b)m}\right]\right\}\delta{\bar{\Gamma}}^{c}{}_{ab}~{}~{}~{}.$ (3.54)
Using Eq. (3.8) we split the variation Eq. 3.53 into two parts
$\delta_{{\bar{\Gamma}}}\mathcal{L}=$ $\delta_{{\hat{\Gamma}}}\mathcal{L}+\delta_{{\tilde{C}}}\mathcal{L}$ (3.55)
$\delta_{{\hat{\Gamma}}}\mathcal{L}=$ $2J_{0}\left\{\tilde{\mathbb{E}}_{c}{}^{ab}+\tfrac{1}{{\rm d}+1}\delta_{c}{}^{(a}\mathbb{E}_{m}{}^{b)m}\right.$ (3.56)
$\qquad\left.-\tfrac{1}{2}{\bar{\nabla}}_{m}\left[\delta_{c}{}^{(a}g^{b)m}\left(\tfrac{1}{4J_{0}\kappa_{0}}+{\mathcal{K}}\right)-4\delta_{c}{}^{(a}{\mathcal{K}}^{b)m}\right]\right\}\delta{\hat{\Gamma}}^{c}{}_{ab}$ (3.57)
$\delta_{{\tilde{C}}}\mathcal{L}=$ $2J_{0}\tilde{\mathbb{E}}_{c}{}^{ab}\delta{\tilde{C}}^{c}{}_{ab}$ (3.58)
where we have exploited the tracelessness of ${\tilde{C}}^{c}{}_{ab}$ to simplify $\delta_{\tilde{C}}\mathcal{L}$ to a single term involving $\tilde{\mathbb{E}}_{c}{}^{ab}$. Thus Eq. (3.58) yields the equations of motion for ${\tilde{C}}^{c}{}_{ab}$ and Eq. (3.56) leads to a contribution to the metric $g^{mn}$ field equations. Summarizing, the equations of motion for ${\tilde{C}}^{c}{}_{ab}$ are
$\tilde{\mathbb{E}}_{c}{}^{ab}\equiv$ $\mathbb{E}_{c}{}^{ab}-\tfrac{1}{{\rm d}+1}\delta_{c}{}^{(a}\mathbb{E}_{m}{}^{b)m}=0$ (3.59)
$\mathbb{E}_{c}{}^{ab}=$${\bar{\nabla}}_{c}\left[g^{ab}\left({\mathcal{K}}+\tfrac{1}{4J_{0}\kappa_{0}}\right)-4{\mathcal{K}}^{ab}\right]-{\bar{\nabla}}_{d}\left(g_{cm}{\mathcal{K}}^{m(ab)d}\right)-\lambda_{0}^{2}{\bar{{\mathcal{P}}}}_{dc}K^{(ab)d}~{}~{}~{}.$ (3.60)
The equations of motion for ${\tilde{C}}^{c}{}_{ab}$ are manifestly covariant and projectively invariant as they are composed entirely of such fields. Where the equations of motion for ${\bar{{\mathcal{P}}}}_{ab}$ and ${\mathcal{D}}_{ab}$ are identical, Eq. (3.59) for ${\tilde{C}}^{c}{}_{ab}$ is not the same as the equations of motion for $\Pi^{c}{}_{ab}$, Eq. (2.63). Utilizing Eq. (3.4) to expand the ¯$\bar{\nabla}_{c}$ derivatives and the definition of ${\bar{{\mathcal{P}}}}_{dc}$, Eq. (3.11), it is straightforward to show that the form of $\mathbb{E}_{c}{}^{ab}$ in Eq. (3.60) is identical to the form in Eq. (2.74). Therefore, the equations of motion for ${\tilde{C}}^{c}{}_{ab}$ indeed are related to those for $\Pi^{c}{}_{ab}$ as in Eq. (2.78). We rewrite the relationship again here in slightly more compact form
$\tilde{\mathbb{E}}_{c}{}^{ab}=$$\tilde{\mathcal{E}}_{c}{}^{ab}-\left(g_{c}\mathbb{D}^{ab}-\tfrac{1}{{\rm d}+1}\delta_{c}{}^{(a}g_{m}\mathbb{D}^{b)m}\right)~{}~{}~{}.$ (3.61)
We see then that the equations of motion for ${\tilde{C}}^{c}{}_{ab}$ and $\Pi^{a}{}_{ab}$ are identical upon enforcing the diff field equation $\mathbb{D}^{ab}=0$.
3.3.3 Field Equations for $g^{mn}$
As as segue to the full equations of motion for $g^{mn}$, we continue with Eq. (3.56) to rewrite it in terms of a variation with respect to the metric $g^{mn}$. We start by contracting the Kronecker delta’s to write this as
$\delta_{{\hat{\Gamma}}}\mathcal{L}=$ $2J_{0}(\tilde{\mathbb{E}}_{c}{}^{ab}\delta{\hat{\Gamma}}^{c}{}_{ab}+V^{a}\delta{\hat{\Gamma}}^{m}{}_{am})$ (3.62)
$V^{a}\equiv$ $\tfrac{2}{{\rm d}+1}\mathbb{E}_{m}{}^{am}-{\bar{\nabla}}_{m}\left[g^{am}\left(\tfrac{1}{4J_{0}\kappa_{0}}+{\mathcal{K}}\right)-4{\mathcal{K}}^{am}\right]$ (3.63)
$=$ $\tfrac{1}{{\rm d}+1}{\bar{\nabla}}_{b}\mathbb{D}^{ab}-\tfrac{\lambda_{0}^{2}}{{\rm d}+1}\left({\bar{\nabla}}_{b}{\bar{\nabla}}_{c}K^{(ab)c}+2{\bar{{\mathcal{P}}}}_{db}K^{bad}\right)$ (3.64)

### v3

@@ III.3 Covariant and Manifestly Projective Invariant TW Field Equations
¯${\bar{\nabla}}_{a}A^{a}=$ $\partial_{a}A^{a}+{\bar{\Gamma}}^{a}{}_{ba}A^{b}$ (140)
$=$ $\partial_{a}A^{a}+\hat{\Gamma}^{a}{}_{ba}A^{b}+{\tilde{C}}^{a}{}_{ba}A^{b}$ (141)
$=$ $\partial_{a}A^{a}+\hat{\Gamma}^{a}{}_{ba}A^{b}$ (142)
$=$ ^$\hat{\nabla}_{a}A^{a}~{}.$ (143)
@@
$\int d^{\rm d}x\sqrt{|g|}\bar{\nabla}_{a}A^{a}=$ $\int d^{\rm d}x\sqrt{|g|}\hat{\nabla}_{a}A^{a}$ (144)
@@
$=$ $\int d^{\rm d}x\partial_{a}(\sqrt{|g|}A^{a})$ (147)
$=$ surface term (148)
where we have used Eq. (36) in going from the second to third lines. This means we can use partial integration with ¯$\bar{\nabla}_{a}$ at the Lagrangian level at no cost to the field equations. For instance given an arbitrary rank three tensor $T^{abc}$
$T^{abc}\bar{\nabla}_{a}\delta\bar{{\mathcal{P}}}_{bc}=$ (¯$-(\bar{\nabla}_{a}T^{abc})\delta\bar{{\mathcal{P}}}_{bc}$ (149)
up to a surface term. Thus, we can work entirely with the TW Lagrangian that we copy from Eqs. (54) and (55) to here for convenience
$\mathcal{L}=$ $-\frac{1}{2\kappa_{0}}(\mathcal{K}+2\Lambda_{0})+\mathbb{L}$ (54)
$\mathbb{L}=$ $J_{0}c\left[\lambda_{0}^{2}K_{bcd}K^{bcd}-\mathcal{K}_{abcd}\mathcal{K}^{abcd}+4\mathcal{K}_{bd}\mathcal{K}^{bd}-\mathcal{K}^{2}\right]$ (55)
and need not work with the full action, Eq. (53).
The connection ¯$\bar{\nabla}_{a}$ provides as much calculational functionality in TW gravity as the Levi-Civita connection ^$\hat{\nabla}_{a}$ does in conventional general relativity. This is a powerful new feature that the projectively invariant and tensorial ¯$\bar{\nabla}_{a}$ and ${\bar{{\mathcal{P}}}}_{ab}$ provide for TW gravity, moving the results of the previous work [20, 21] closer to the language of more conventional language in general relativity. In the following, we will demonstrate the utility of this new calculation tool.
III.3.1 Field Equations for ${\bar{{\mathcal{P}}}}_{ab}$
@@
$\delta_{{\bar{{\mathcal{P}}}}}{\mathcal{K}}^{a}{}_{bcd}=$ $\delta^{a}{}_{[c}\delta{\bar{{\mathcal{P}}}}_{d]b}~{}~{}~{},~{}~{}~{}\delta_{{\bar{{\mathcal{P}}}}}K_{bcd}={\bar{\nabla}}_{[c}\delta{\bar{{\mathcal{P}}}}_{d]b}$ (150)
@@
$\delta_{{\bar{{\mathcal{P}}}}}\mathcal{L}=-\tfrac{1}{2\kappa_{0}}g^{ab}\delta_{{\bar{{\mathcal{P}}}}}{\mathcal{K}}_{ab}+2J_{0}\lambda_{0}^{2}K^{bcd}\delta_{{\bar{{\mathcal{P}}}}}K_{bcd}$ (151)
$\qquad\quad-2J_{0}{\mathcal{K}}_{a}^{\ bcd}\delta_{{\bar{{\mathcal{P}}}}}{\mathcal{K}}^{a}{}_{bcd}-2J_{0}(-4{\mathcal{K}}^{ab}+{\mathcal{K}}g^{ab})\delta_{{\bar{{\mathcal{P}}}}}{\mathcal{K}}_{ab}$ (152)
$=-\tfrac{1}{2\kappa_{0}}g^{ab}({\rm d}-1)\delta{\bar{{\mathcal{P}}}}_{ab}+2J_{0}\lambda_{0}^{2}K^{bcd}{\bar{\nabla}}_{[c}\delta{\bar{{\mathcal{P}}}}_{d]b}$ (153)
$-2J_{0}{\mathcal{K}}_{a}^{\ bcd}\delta^{a}_{\ [c}\delta{\bar{{\mathcal{P}}}}_{d]b}-2J_{0}(-4{\mathcal{K}}^{ab}+{\mathcal{K}}g^{ab})({\rm d}-1)\delta{\bar{{\mathcal{P}}}}_{ab}~{}~{}~{}~{}~{}~{}$ (154)
where recall here on $\mathcal{M}$ indices are raised and lowered via $g^{ma}$ and $g_{am}$ as in Eqs. (57), (58), and (59). Integrating by parts in the second term, contracting indices in the third, and renaming and reordering indices yields
$\delta_{{\bar{{\mathcal{P}}}}}\mathcal{L}=$ $\big{(}-\tfrac{{\rm d}-1}{2\kappa_{0}}g^{ab}+2J_{0}\lambda_{0}^{2}{\bar{\nabla}}_{c}K^{(ab)c}-2J_{0}({\rm d}-1){\mathcal{K}}g^{ab}$ (155)
$\quad+8J_{0}({\rm d}-1){\mathcal{K}}^{ab}-4J_{0}{\mathcal{K}}^{ab}\big{)}\delta{\bar{{\mathcal{P}}}}_{ab}$ (156)
where we have used the antisymmetry of the last two indices of $K^{abc}$ to reorder the indices.
Simplifying we have
$\delta_{{\bar{{\mathcal{P}}}}}\mathcal{L}=$ $2J_{0}\mathbb{D}^{ab}\delta{\bar{{\mathcal{P}}}}_{ab}$ (157)
@@
$0=\mathbb{D}^{ab}=$ $\lambda_{0}^{2}\bar{\nabla}_{c}K^{(ab)c}-({\rm d}-1)\mathbb{K}^{ab}-2\mathcal{K}^{ab}~{}.$ (158)
Recall the projectively invariant tensor $\mathbb{K}^{ab}$ was defined in Eq. (82). Using Eq. (106) to expand the first term, it is straightforward to show that this is the exact same as Eq. (81) that was derived via variation with respect to ${\mathcal{D}}_{ab}$. In the form of Eq. (158), however, we see this equation of motion is covariant and manifestly projectively invariant as it is written entirely in terms of such objects.
III.3.2 Field Equations for ${\tilde{C}}^{c}{}_{ab}$
@@
$\delta_{{\bar{\Gamma}}}{\mathcal{K}}^{a}{}_{bcd}=$${\bar{\nabla}}_{[c}\delta{\bar{\Gamma}}^{a}{}_{d]b}~{}~{}~{},~{}~{}~{}\delta_{{\bar{\Gamma}}}K_{bcd}=-\delta{\bar{\Gamma}}^{m}{}_{b[c}{\bar{{\mathcal{P}}}}_{d]m}$ (159)
@@
$\delta_{{\bar{\Gamma}}}\mathcal{L}=-\tfrac{1}{2\kappa_{0}}g^{ab}\delta_{{\bar{\Gamma}}}{\mathcal{K}}_{ab}+2J_{0}\lambda_{0}^{2}K^{bcd}\delta_{{\bar{\Gamma}}}K_{bcd}$ (160)
$\qquad\quad-2J_{0}{\mathcal{K}}_{a}^{\ bcd}\delta_{{\bar{\Gamma}}}{\mathcal{K}}^{a}{}_{bcd}-2J_{0}(-4{\mathcal{K}}^{ab}+{\mathcal{K}}g^{ab})\delta_{{\bar{\Gamma}}}{\mathcal{K}}_{ab}$ (161)
$=-\tfrac{1}{2\kappa_{0}}g^{ab}{\bar{\nabla}}_{[c}\delta{\bar{\Gamma}}^{c}{}_{b]a}+2J_{0}\lambda_{0}^{2}K^{bcd}(-\delta{\bar{\Gamma}}^{m}{}_{b[c}{\bar{{\mathcal{P}}}}_{d]m})$ (162)
$-2J_{0}{\mathcal{K}}_{a}^{\ bcd}{\bar{\nabla}}_{[c}\delta{\bar{\Gamma}}^{a}_{\ d]b}-2J_{0}(-4{\mathcal{K}}^{ab}+{\mathcal{K}}g^{ab}){\bar{\nabla}}_{[c}\delta{\bar{\Gamma}}^{c}{}_{b]a}$ (163)
@@
where recall here on $\mathcal{M}$ indices are raised and lowered via $g^{ma}$ and $g_{am}$ as in Eqs. (57), (58), and (59). Integrating by parts, relabeling some indices, and collecting like terms, such as the first and last terms that form a $\mathbb{K}^{ab}$ term (Eq. (82)), and simplifying results in
$\delta_{{\bar{\Gamma}}}\mathcal{L}=2J_{0}\delta{\bar{\Gamma}}^{c}_{\ a[b}{\bar{\nabla}}_{c]}\mathbb{K}^{ab}$ (164)
@@
$\qquad+2J_{0}\left[-\lambda_{0}^{2}K^{(ab)d}{\bar{{\mathcal{P}}}}_{dc}+{\bar{\nabla}}_{d}{\mathcal{K}}_{c}^{\ (b|d|a)}\right]\delta{\bar{\Gamma}}^{c}_{\ ab}~{}~{}~{}~{}~{}~{}~{}$ (165)
@@
where vertical lines denote exclusion from the symmetrization parenthesis. Introducing a Kronecker delta symbols in the first term to move the antisymmetrization of indices off of $\delta{\bar{\Gamma}}^{c}{}_{ab}$ and ¯${\bar{\nabla}}_{c}$, using the antisymmetry property ${\mathcal{K}}^{mbda}=-{\mathcal{K}}^{mbad}$ in the second term on the second line, and simplifying results in
$\delta_{{\bar{\Gamma}}}\mathcal{L}=2J_{0}[\delta^{a}_{\ [m}{\bar{\nabla}}_{c]}\mathbb{K}^{bm}-\lambda_{0}^{2}{\bar{{\mathcal{P}}}}_{dc}K^{(ab)d}-{\bar{\nabla}}_{d}{\mathcal{K}}_{c}^{\ (ab)d}]\delta{\bar{\Gamma}}^{c}_{\ ab}.$ (166)
As a quick aside, we note that the tensor $\mathbb{E}_{c}{}^{ab}$ defined in Eq. (88) can be expressed as
$\mathbb{E}_{c}{}^{ab}=$${\bar{\nabla}}_{c}\mathbb{K}^{ab}-{\bar{\nabla}}_{d}{\mathcal{K}}_{c}^{\ (ab)d}-\lambda_{0}^{2}{\bar{{\mathcal{P}}}}_{dc}K^{(ab)d}~{}.$ (167)
Utilizing Eq. (106) to expand the ¯$\bar{\nabla}_{c}$ derivatives and with the definition of ${\bar{{\mathcal{P}}}}_{dc}$, Eq. (112), it is straightforward to show that Eq. (167) indeed reduces to Eq. (88). With our newfound expression for $\mathbb{E}_{c}{}^{ab}$ in Eq. (167), the variation Eq. (166) can be cast in the following form
$\delta_{{\bar{\Gamma}}}\mathcal{L}=$ $2J_{0}\Big{[}\mathbb{E}_{c}{}^{ab}-\delta^{a}_{\ c}{\bar{\nabla}}_{m}\mathbb{K}^{bm}\Big{]}\delta{\bar{\Gamma}}^{c}_{\ ab}~{}.$ (168)
@@
To facilitate the goal of simultaneously isolating the equations of motion for ${\tilde{C}}^{c}{}_{ab}$ whilst keeping track of the contributions to the equations of motion for the metric $g^{mn}$, we substitute in the traceless combination $\tilde{\mathbb{E}}_{c}{}^{ab}$ defined in Eq. (93) to express Eq. (168) as
@@
$\delta_{{\bar{\Gamma}}}\mathcal{L}=2J_{0}\left\{\tilde{\mathbb{E}}_{c}^{\ ab}+\tfrac{2}{{\rm d}+1}\delta_{c}^{\ a}\mathbb{E}_{m}^{\ bm}-\delta_{c}^{\ a}{\bar{\nabla}}_{m}\mathbb{K}^{bm}\right\}\delta{\bar{\Gamma}}^{c}_{\ ab}.$ (169)
Collecting the rightmost two terms, this expression simplifies to
$\delta_{{\bar{\Gamma}}}\mathcal{L}=$ $2J_{0}(\tilde{\mathbb{E}}_{c}{}^{ab}\delta{\bar{\Gamma}}^{c}{}_{ab}+V^{a}\delta{\bar{\Gamma}}^{m}{}_{am})$ (170)
$V^{a}\equiv$ $\tfrac{2}{{\rm d}+1}\mathbb{E}_{m}{}^{am}-{\bar{\nabla}}_{m}\mathbb{K}^{am}$ (171)
$=$ $\tfrac{1}{{\rm d}+1}{\bar{\nabla}}_{b}\mathbb{D}^{ab}-\tfrac{\lambda_{0}^{2}}{{\rm d}+1}\left({\bar{\nabla}}_{b}{\bar{\nabla}}_{c}K^{(ab)c}+2{\bar{{\mathcal{P}}}}_{db}K^{bad}\right)~{}~{}~{}~{}~{}~{}$ (172)